OpenStudy (anonymous):

The English Alphabet has 20 letters more than one-fourth the number of letters in the Greek Alphabet. The English Alphabet has 26 letters. Which of the following equations would you use to find the number of letters, g, in the Greek Alphabet? (5 points) The sum of g and 20 all over 4 equals 26 4g − 20 = 26 1 over 4 g + 20 = 26 4g − 26 = 20

Directrix (directrix):

The English Alphabet has 20 letters more than one-fourth the number of Greek letters 26 = 20 + (1/4)* g @hallway Does that equation make sense? Do you think it is correct so far?

OpenStudy (anonymous):

yeah that is sort of what I got, but the options don't make sense for me

Directrix (directrix):

Isn't what I wrote : 26 = 20 + (1/4)* g the same as the option: 1 over 4 g + 20 = 26

Directrix (directrix):

26 = 20 + (1/4)* g 20 + (1/4)* g = 26 (1/4) g + 20 = 26 one-fourth g + 20 = 26